Calculate the number of grams of oxygen required to convert 43.0 g of glucose to CO2, and H2O

Glucose, C6H12O6, is used as an energy source by the human body. The overall reaction in the body is described by the equation

C6H12O6(aq) + 6O2(g) — 6CO2(g) + 6H2O(1) 

Calculate the number of grams of oxygen required to convert 43.0 g of glucose to CO2, and H2O. 

mass of O2 = _______ g 

Calculate the number of grams of CO2, produced. 

mass of CO2 = _______ g 
Answer #1

1)

Molar mass of C6H12O6,

MM = 6*MM(C) + 12*MM(H) + 6*MM(O)

= 6*12.01 + 12*1.008 + 6*16.0

= 180.156 g/mol

mass of C6H12O6 = 43 g

mol of C6H12O6 = (mass)/(molar mass)

= 43/1.802*10^2

= 0.2387 mol

According to balanced equation

mol of O2 required = (6/1)* moles of C6H12O6

= (6/1)*0.2387

= 1.432 mol

Molar mass of O2 = 32 g/mol

mass of O2 = number of mol * molar mass

= 1.432*32

= 45.83 g

Answer: 45.8 g

2)

According to balanced equation

mol of CO2 formed = (6/1)* moles of C6H12O6

= (6/1)*0.2387

= 1.432 mol

Molar mass of CO2,

MM = 1*MM(C) + 2*MM(O)

= 1*12.01 + 2*16.0

= 44.01 g/mol

mass of CO2 = number of mol * molar mass

= 1.432*44.01

= 63.03 g

Answer: 63.0 g
Answer #2

C6H12O6 (aq) + 69 (9) 6co(g) + 6H₂O (1) Molecular weight of Hyz 05 = 180 gm./mol Molecular weight of Oz -32 gmi mol Molecular= 45. 84 8m @ Mass of o, = 45 - 81 эт. 180 gm. c6 Hz of reacts with (3246)gmę nd produce (44x6) gm co , т а т с 2 б ю:u pro d